OpenStudy (anonymous):

Solve 7(x - 4) + 4x = 71. 11 9 7 5

OpenStudy (anonymous):

do you know how to distribute?

OpenStudy (nurali):

7(x - 4) + 4x = 71. 7x-28+4x=71 11x=71+28 11x=99 x=9

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

yes